### v1

@@ 6.1 Formal Proofs of Quantum Forgetting
Classical guarantees rely on retraining equivalence, influence corrections, or stability bounds. Quantum settings require a different contract: suppression of retrievable information within CPTP maps. A useful target is to certify that fidelity and trace distance cross pre-set margins between pre- and post-unlearning states, that is $F(\rho,\rho^{\prime})\leq\epsilon$ and $D_{\mathrm{tr}}(\rho,\rho^{\prime})\geq\delta$, for task-specific thresholds [126, 127, 128]. These conditions pair naturally with the data-processing inequality and turn hardware noise into conservative evidence rather than a nuisance.
Parameterized quantum circuits offer structure for proofs. Modular architectures allow selective resets or stochastic reinitializations of forgetting blocks that isolate the contribution of marked data or clients [126]. After a reset, vanishing local gradients $\partial_{\theta_{i}}\mathcal{L}=0$ signal orthogonality to forgotten modes and support geometric certificates. Architectural patterns strengthen this path. Decoherence-mediated forgetting uses controlled channels $\mathcal{E}_{p}$ to raise entropy on targeted subspaces. Variational reset reinitializes blocks and bounds post-tuning drift. Hybrid modularity separates quantum edits from classical feedback to prevent re-entanglement [127]. Bridging practice and proof will also need empirical-to-formal pipelines. Hybrid models that couple classical encoders with variational circuits retain accuracy after local resets and supply testbeds for certificate design [129]. Verification tools such as VeriQR, designed for robustness, can reason about reachability before and after unlearning and check disjointness of state sets within a contractive metric [128]. These steps align theory with thermodynamics and move QMU toward machine-checkable guarantees.
@@ 6.2 Scalability and Secure Architectures
Scalability is a joint hardware–systems problem. NISQ limits on depth and connectivity require modular subsystems with classical orchestration and explicit verifiability. Reservoir computing offers a principled scaffold. Classical and quantum reservoirs are universal for fading-memory functionals, so local readout resets can approximate global retraining with bounded error [130]. This supports distributed unlearning without full model restart.
Fully quantum generative models are another lever. QGANQB keeps both generator and discriminator quantum, enabling native process-level unlearning and avoiding classical bottlenecks [131]. Quantum imaging adds a complementary knob. Compact encodings such as FRQI and NEQR permit localized spectral edits (Fourier or Haar) that implement class- or region-specific forgetting with polynomial resources [132]. Security must be baked into these designs. Controlled entanglement interfaces and modular middleware isolate subsystems and limit cross-node leakage in federated settings [127]. This architectural security complements QDP and homomorphic transport and supports the secure-composition guarantees needed for QFU at scale.
@@ 6.3 Post-Unlearning Interpretability
Erasure should not obscure why the model still works. Interpretability after unlearning is an information allocation problem: reduce $I(\rho;\rho^{\prime})$ while preserving explanatory power under relevant observables. Self-adaptive quantum kernel PCA can re-optimize components to suppress variance tied to removed samples and keep embeddings faithful to the retained subspace [133]. Physics-informed hybrids, such as TE-QPINNs, embed constraints from governing equations so that deletion does not violate domain knowledge [134]. Architectural lineage also helps. Evolutionary records in hybrids like HQCNN-REGA trace which blocks were altered and support audit-ready explanations [135]. Formally, one can target $\min I(\rho;\rho^{\prime})$ subject to $\mathcal{I}(\rho^{\prime},\mathcal{O})\geq\tau$ for chosen observables $\mathcal{O}$ and threshold $\tau$. This framing matches the sensitivity tools in Sec. 4 and yields interpretable, deletion-aware reports.
@@ 6.4 Ethics and Societal Implications
QMU must earn trust through verification and fairness. Verifiable delegation allows clients to check remote computations and supports auditable unlearning on untrusted hardware [136]. Biomedical deployments underscore the need for human oversight and reproducibility when QSVCs approach classical accuracy [137]. Hardware fairness is a new axis. Noise varies by device, which can induce unequal reliability. Calibration transparency, cross-device reproducibility, and bounds on trace-distance divergence should therefore be part of governance checklists [138]. In federated settings, deletion proofs and DP accounting must be first-class artifacts, not afterthoughts. These practices connect directly to the metrics in Sec. 5.1 and prepare the ecosystem for standardized audits.
@@
Provable forgetting Modular PQCs; CPTP-based certificates; VeriQR-style reachability [126, 127, 128] Certified bounds on $F$ and $D_{\mathrm{tr}}$; machine-checkable proofs.
Scalable QML/QFU Reservoir universality; fully quantum GANs; secure modular middleware [130, 131, 127] Sub-linear retraining cost; elastic depth/width; efficient client-level deletion.
Post-unlearning interpretability SAQK-PCA; TE-QPINNs; evolutionary lineage (HQCNN-REGA) [133, 134, 135] Observable-preserving edits; quantitative interpretability indices; audit trails.
Ethical governance Verifiable delegation; calibration fairness; reproducible diagnostics [136, 138, 137] Standardized reports; backend-agnostic audits; accountable QFU pipelines.

### v2

@@ 6.1 Formal Proofs of Quantum Forgetting
Classical guarantees rely on retraining equivalence, influence corrections, or stability bounds. Quantum settings require a different contract: suppression of retrievable information within CPTP maps. A useful target is to certify that fidelity and trace distance cross pre-set margins between pre- and post-unlearning states, that is $F(\rho,\rho^{\prime})\leq\epsilon$ and $D_{\mathrm{tr}}(\rho,\rho^{\prime})\geq\delta$, for task-specific thresholds [nhlapo2025zerosvm, klymenko2025architecturalpatterns, lin2025veriqr]. These conditions pair naturally with the data-processing inequality and turn hardware noise into conservative evidence rather than a nuisance.
Parameterized quantum circuits offer structure for proofs. Modular architectures allow selective resets or stochastic reinitializations of forgetting blocks that isolate the contribution of marked data or clients [nhlapo2025zerosvm]. After a reset, vanishing local gradients $\partial_{\theta_{i}}\mathcal{L}=0$ signal orthogonality to forgotten modes and support geometric certificates. Architectural patterns strengthen this path. Decoherence-mediated forgetting uses controlled channels $\mathcal{E}_{p}$ to raise entropy on targeted subspaces. Variational reset reinitializes blocks and bounds post-tuning drift. Hybrid modularity separates quantum edits from classical feedback to prevent re-entanglement [klymenko2025architecturalpatterns]. Bridging practice and proof will also need empirical-to-formal pipelines. Hybrid models that couple classical encoders with variational circuits retain accuracy after local resets and supply testbeds for certificate design [guha2025resq]. Verification tools such as VeriQR, designed for robustness, can reason about reachability before and after unlearning and check disjointness of state sets within a contractive metric [lin2025veriqr]. These steps align theory with thermodynamics and move QMU toward machine-checkable guarantees.
@@ 6.2 Scalability and Secure Architectures
Scalability is a joint hardware–systems problem. NISQ limits on depth and connectivity require modular subsystems with classical orchestration and explicit verifiability. Reservoir computing offers a principled scaffold. Classical and quantum reservoirs are universal for fading-memory functionals, so local readout resets can approximate global retraining with bounded error [monzani2025universality]. This supports distributed unlearning without full model restart.
Fully quantum generative models are another lever. QGANQB keeps both generator and discriminator quantum, enabling native process-level unlearning and avoiding classical bottlenecks [ding2025qganqb]. Quantum imaging adds a complementary knob. Compact encodings such as FRQI and NEQR permit localized spectral edits (Fourier or Haar) that implement class- or region-specific forgetting with polynomial resources [farooq2025quantumimage]. Security must be baked into these designs. Controlled entanglement interfaces and modular middleware isolate subsystems and limit cross-node leakage in federated settings [klymenko2025architecturalpatterns]. This architectural security complements QDP and homomorphic transport and supports the secure-composition guarantees needed for QFU at scale.
@@ 6.3 Post-Unlearning Interpretability
Erasure should not obscure why the model still works. Interpretability after unlearning is an information allocation problem: reduce $I(\rho;\rho^{\prime})$ while preserving explanatory power under relevant observables. Self-adaptive quantum kernel PCA can re-optimize components to suppress variance tied to removed samples and keep embeddings faithful to the retained subspace [wang2025saqkpca]. Physics-informed hybrids, such as TE-QPINNs, embed constraints from governing equations so that deletion does not violate domain knowledge [berger2025teqpin]. Architectural lineage also helps. Evolutionary records in hybrids like HQCNN-REGA trace which blocks were altered and support audit-ready explanations [mukhanbet2025hqcnnrega]. Formally, one can target $\min I(\rho;\rho^{\prime})$ subject to $\mathcal{I}(\rho^{\prime},\mathcal{O})\geq\tau$ for chosen observables $\mathcal{O}$ and threshold $\tau$. This framing matches the sensitivity tools in Sec. 4 and yields interpretable, deletion-aware reports.
@@ 6.4 Ethics and Societal Implications
QMU must earn trust through verification and fairness. Verifiable delegation allows clients to check remote computations and supports auditable unlearning on untrusted hardware [yang2025verifiablevqa]. Biomedical deployments underscore the need for human oversight and reproducibility when QSVCs approach classical accuracy [passo2025ibmbrain]. Hardware fairness is a new axis. Noise varies by device, which can induce unequal reliability. Calibration transparency, cross-device reproducibility, and bounds on trace-distance divergence should therefore be part of governance checklists [wang2025noisykernel]. In federated settings, deletion proofs and DP accounting must be first-class artifacts, not afterthoughts. These practices connect directly to the metrics in Sec. 5.1 and prepare the ecosystem for standardized audits.
@@
Provable forgetting Modular PQCs; CPTP-based certificates; VeriQR-style reachability [nhlapo2025zerosvm, klymenko2025architecturalpatterns, lin2025veriqr] Certified bounds on $F$ and $D_{\mathrm{tr}}$; machine-checkable proofs.
Scalable QML/QFU Reservoir universality; fully quantum GANs; secure modular middleware [monzani2025universality, ding2025qganqb, klymenko2025architecturalpatterns] Sub-linear retraining cost; elastic depth/width; efficient client-level deletion.
Post-unlearning interpretability SAQK-PCA; TE-QPINNs; evolutionary lineage (HQCNN-REGA) [wang2025saqkpca, berger2025teqpin, mukhanbet2025hqcnnrega] Observable-preserving edits; quantitative interpretability indices; audit trails.
Ethical governance Verifiable delegation; calibration fairness; reproducible diagnostics [yang2025verifiablevqa, wang2025noisykernel, passo2025ibmbrain] Standardized reports; backend-agnostic audits; accountable QFU pipelines.